A light source emits a beam of photons, each of which has a momentum of 3.00 × 10 − 29 kg.m/s. What is the frequency of the photons?
Answer #1

P= E/c=hfc

Frequency is given by

f = Pc/h

= (3×10-29) ×(3×108)/(6.626×10-34)

= 1.36 × 1013 Hz